OpenStudy (anonymous):

PLZ SOMEONE HELP IVE BEEN WORKING ON THIS ONE QUESTION FOR 3 HOURS 3 FREAKING OURS SO SOMEONE PLZ HELP ILL GIVE MEDAL

OpenStudy (anonymous):

Whats the question?

OpenStudy (pixiedust1):

Sure buddy! I'm having one of those days too... hopefully we can help :)

OpenStudy (anonymous):

THANK YOU ALL BLESS THE LORD

OpenStudy (dbzfan836):

Im also havin a hard time[derp]

OpenStudy (pixiedust1):

LOL! Confused people... UNITE!

OpenStudy (anonymous):

AND I HAVE ANOTHER TEST AFTER THIS im having a bad day hold on iget estion

OpenStudy (anonymous):

@Dbzfan836 tag me in ill see if i can help

OpenStudy (anonymous):

Write a word problem that can be solved by working backward. Include at least two steps in the problem. Then, show the answer and how the problem can be solved by working backward.

OpenStudy (dbzfan836):

b kjbjkcvbdfkcvjerkvnjnsdcvkmasnd fvkjrnc;vhjbvkjbrvkjbsdfvfs is my way of saying i feel bad

OpenStudy (anonymous):

ha thx

OpenStudy (anonymous):

Carli filled her tank full with gas. She used half her gas to go to school, then 5 gallons going to work, and the last 3 gallons just driving around. How much can her tank hold? answer is 16 gallons

OpenStudy (anonymous):

all you guys are awesome

OpenStudy (pixiedust1):

Nice one cheif!

OpenStudy (pixiedust1):

Any other questions @i_need_help123

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

chief it cant be copied has to be own words no copy write

OpenStudy (anonymous):

Yup just reword it easy piser lemon squesie

OpenStudy (anonymous):

I cant

OpenStudy (anonymous):

lol pise i ment

OpenStudy (anonymous):

my teacher looks all the stuff up

OpenStudy (pixiedust1):

Or just switch the subject. Instead of gas, make it candy. She ate ____ on the way to school, ___ on the way to work __ and ___ when just driving around

OpenStudy (pixiedust1):

Get what I mean?

OpenStudy (anonymous):

nice @PixieDust1

OpenStudy (anonymous):

ok but plz plz dnt leave

OpenStudy (pixiedust1):

thanks bro

OpenStudy (pixiedust1):

And this is an online tutoring site, not a cheating site. We're da tutors! :P

OpenStudy (anonymous):

we wont i gotta just finish my Test in apex ill keep this tab open tag me in it if you need help

OpenStudy (pixiedust1):

jk lol

OpenStudy (pixiedust1):

Good luck @Cheifkeffer

OpenStudy (anonymous):

lol ya i need it this dam geo is hard but im almost done with this semester @PixieDust1

OpenStudy (pixiedust1):

Aww lol And @Cheifkeffer that's a lot of money you got in your profile pic lol! :D

OpenStudy (anonymous):

ok im back

OpenStudy (pixiedust1):

Okay! Hello! :)

OpenStudy (anonymous):

@PixieDust1 and @Cheifkeffer

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

can you help wth some more

OpenStudy (pixiedust1):

Yuppers! Present and ready for action

OpenStudy (anonymous):

on this one I have to get a 10 percent

OpenStudy (anonymous):

100%

OpenStudy (pixiedust1):

Pressure lol

OpenStudy (pixiedust1):

I'm gonna tag a mod to check our work then

OpenStudy (anonymous):

In your own words, explain two ways to determine if an equation is linear.

OpenStudy (anonymous):

if it can be put in the standard line formula of y = mx + b if it can be put in the standard slope formula of (y2 - y1)/(x2 - x1) = m

OpenStudy (pixiedust1):

1.) No idea 2.) No idea Lol! I'm gonna tag a mod, if you don't mind :D

OpenStudy (anonymous):

@i_need_help123

OpenStudy (pixiedust1):

YAY CHEIF!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

k give me a sec

OpenStudy (pixiedust1):

You back?

OpenStudy (anonymous):

im back

OpenStudy (anonymous):

my comp. froze

OpenStudy (anonymous):

Which of the following equations is graphed as a straight line? y equals 3 x y equals 4 divided by x y equals 7 x squared x y equals 8

OpenStudy (anonymous):

@Cheifkeffer @PixieDust1

OpenStudy (anonymous):

help

OpenStudy (perl):

bs, i dont believe you spent 3 hours on this

OpenStudy (anonymous):

is it multiple choice

OpenStudy (anonymous):

@i_need_help123

OpenStudy (anonymous):

y equals 3 x

OpenStudy (anonymous):

Fun Lanes Bowling charges a fee of $4 to use their shoes and $6 for each game played. Salim wants to figure out how much it will cost him to bowl. Part 1: Write an equation representing this problem. Let x be the number of games bowled, and let y be the total cost. Part 2: Is x the dependent variable or the independent variable? Using complete sentences, explain how you know.

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

part 1 is : 4 + 6x = y becuase 4 is a given, because you only pay once to use the shoes. To get the total cost of the number of games bowled, multiply 6 by the number of games. 6x Add the two together to get the total cost. part 2: For part 2, x is an independent variable because its value can be changed without making the equation untrue

OpenStudy (pixiedust1):

4x+6=y

OpenStudy (anonymous):

member my teacher looks everything up

OpenStudy (anonymous):

i have been typing for a lil haha

OpenStudy (anonymous):

no she looks up the question

OpenStudy (pixiedust1):

Haha lol

OpenStudy (anonymous):

wen she is grading

OpenStudy (anonymous):

what a mean teacher :(

OpenStudy (anonymous):

so chief did u look it up

OpenStudy (pixiedust1):

Look what up?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

the answer

OpenStudy (anonymous):

If you are scared she might think someone gave it to you just re word it

OpenStudy (anonymous):

uh did u im not trying to be mean but this happened to me before

OpenStudy (anonymous):

soooo

OpenStudy (anonymous):

canu help wth one more this ia hard one

OpenStudy (anonymous):

Here is a table of values for the linear equation y equals 2 x minus 4. The x value is 3. The y value is missing. What is the missing number? nineteen twenty-four ten 2

OpenStudy (anonymous):

@PixieDust1

OpenStudy (anonymous):

2

OpenStudy (pixiedust1):

y=2x-4 let's put in 3 for x y=2*3-4 2*3=6 6-4=2 Therefore, y=2

OpenStudy (anonymous):

@PixieDust1 nice

OpenStudy (pixiedust1):

Thank ya! Those kinds of problems are so much FUN!! :D

OpenStudy (pixiedust1):

@i_need_help123 Does that help?